# Incentive Compatibility for AI Alignment in Sociotechnical Systems: Positions and Prospects

### v1

@@ 1 Introduction
“If we use, to achieve our purposes, a mechanical agency with whose operation we cannot interfere effectively $\dots$ we had better be quite sure that the purpose put into the machine is the purpose which we really desire.” (Wiener, 1960; Russell, 2019)
Norbert Wiener
@@
The rapid development of artificial intelligence (AI) has had a significant impact on human society (Makridakis, 2017; Peeters et al., 2021; Wamba et al., 2021), from robots entering human production and living environments (Michaelis & Mutlu, 2018; Fu et al., 2024) to large language models (LLMs) capable of complex natural language interactions (Zhao et al., 2023; Bubeck et al., 2023) and reasoning ability (Wei et al., 2022; Wang et al., 2022). Therefore, an increasing number of people believe that as AI capabilities improve, AI systems will become integrated into human society in the future and be deployed in increasingly complex scenarios (Gladden, 2019; Dwivedi et al., 2021). Conversely, the powerful capabilities of AI systems have raised concerns about their safety (Cath et al., 2018; Peeters et al., 2021), especially considering their behavioral motivations  [^1] https://www.scai.gov.sg/scai-question-6/, alignment science  [^2] https://www.anthropic.com/news/core-views-on-ai-safety and how they align with human values and intentions (Ji et al., 2023). This is recognized as the “AI Alignment” problem.
@@
Substantial progress has been made in addressing AI alignment issues, especially in the forward alignment process (Ji et al., 2023), which enables AI systems to have alignment capabilities 2. The methods for this process can mainly be divided into two categories. The first category involves learning from feedback (Christiano et al., 2017; Ziegler et al., 2019; Bai et al., 2022a; Ouyang et al., 2022), and there have been some significant research topics, including preference modeling (Akrour et al., 2011; Wirth et al., 2017), policy learning (Amodei et al., 2016; Ibarz et al., 2018), and scalable oversight (Christiano et al., 2018; Irving et al., 2018; Bai et al., 2022a; Burns et al., 2023). The second category focuses on resolving distributional shift (Di Langosco et al., 2022; Ngo et al., 2022) in learning, with notable subproblems including algorithmic interventions (Vapnik, 1991; Krueger et al., 2021; Lubana et al., 2023), adversarial training (Goodfellow et al., 2014; Poursaeed et al., 2021), and cooperative training (Dafoe et al., 2020, 2021).
@@
However, these methods only consider the given alignment objectives, focusing solely on technical components such as dataset, architecture, and training algorithms, etc. (Weidinger et al., 2023b), overlooking the fact that AI systems are sociotechnical systems (Selbst et al., 2019). Some studies have indicated that relying solely on technical means will result in a sociotechnical gap between the model’s development context and its actual deployment context (Selbst et al., 2019; Lazar & Nelson, 2023), which is also detrimental to further social governance. Such examples are not uncommon in daily life. ChatGPT, trained on internet data and fine-tuned through RLHF (Ouyang et al., 2022), requires prompt engineering for adaptation to individual needs, highlighting unaddressed challenges in existing alignment techniques. Additionally, for sociotechnical systems, existing research is more concerned with only societal components like governance and evaluation methods (Raji et al., 2020; Dean et al., 2021; Weidinger et al., 2023b). Thus, currently, there are lack of means to simultaneously consider both technical and societal components, enabling AI systems themselves to maintain consensus with human society.
@@
Incentive Compatibility (IC) (Hurwicz, 1972), derived from game theory, suggests that participants only need to pursue their true interests to reach optimal outcomes (Roughgarden, 2010). This concept leverages self-interested behavior, aligning actions with the game designer’s goals (Groves & Ledyard, 1987). With IC, each agent can maintain private goal information acquired during pretraining. Only by reconstructing different environments and rules,agents can optimize their own objectives to achieve outcomes that meet the needs of human society in different contexts. Therefore, we believe that exploring the IC property for AI alignment problems in sociotechnical systems is a highly worthwhile research endeavor.
In this paper, we separate a new subproblem from AI alignment problems in sociotechnical systems, called Incentive Compatibility Sociotechnical Alignment Problem (ICSAP), and based on this, we propose our main position:
@@
Achieving incentive compatibility can simultaneously consider both technical and societal components in the forward alignment phase, enabling AI systems to keep consensus with human societies in different contexts.
@@ 2 Motivation and Opportunity: A Brief Example
In this section, we will use a very simple example to demonstrate how IC works in addressing AI alignment issues in sociotechnical systems through mechanism design which will be illustrated specifically in section 5.
@@
Consider a classic divide and choose problem: the two-player cake cutting (Steinhaus, 1948). In this example (see Figure 1), two self-interested agents aim to maximize their cake share, while the human seeks an equal division. If either agent cuts the cake, they’ll take the whole. To align individual interests with the human’s goal, a simple mechanism is proposed: the cutter chooses second. This constraint ensures the agent’s pursuit of self-interest coincides with the host’s objective, achieving Alignment. The mechanism’s IC conditions facilitate consensus on an equal distribution and maximization of cake consumption.
If we consider the agent as an AI system and its desire to eat the most cake as the objective imparted by the technical component of training, we only need to use automated methods to search for corresponding mechanisms with IC properties as rules based on different real-world requirements to effectively solve ICSAP. Of course, hosts can also have different contextual needs, and they may not necessarily be self-interested. Here, we’re just using this one example to illustrate our point.
3 Related Work
To better understand the relevant content of ICSAP, in this section, we will introduce background work from two parts: AI alignment and sociotechnical systems.
AI Alignment.
The scope of AI alignment issues can mainly be divided into two processes, namely forward alignment and backward alignment (Ji et al., 2023).
The former refers to the development training and deployment training stages aimed at endowing AI systems with alignment capabilities. This process has yielded many significant achievements, including RLHF (Christiano et al., 2017; Ziegler et al., 2019; Bai et al., 2022a; Ouyang et al., 2022), DPO (Rafailov et al., 2023), etc., which align LLMs through fine-tuning methods; RLAIF (Bai et al., 2022b) RLHAIF (Cotra, 2021) IDA (Christiano et al., 2018) RRM (Leike et al., 2018; Hubinger, 2020) Debate (Irving et al., 2018; Michael et al., 2023) CIRL (Hadfield-Menell et al., 2016) Weak-To-Strong (Burns et al., 2023) and other scalable oversight methods; as well as REx (Krueger et al., 2021) CBFT (Lubana et al., 2023) and other works addressing distribution shift issues.
The latter mainly refers to the process of ensuring that AI systems can meet alignment-related requirements, including the assurance and governance phases. Current work includes ETHICS (Hendrycks et al., 2020), SOCIAL-CHEM-101 (Forbes et al., 2020), MIC (Ziems et al., 2022), TRUSTLLM (Sun et al., 2024), etc., which involve safety evaluation datasets. Methods for handling failure modes through adversarial attacks via red teaming (Perez et al., 2022; Liu et al., 2023); research on mechanistic interpretability (Olah et al., 2020) and governance criteria like ARC Evals  [^3] https://metr.org/blog/2023-03-18-update-on-recent-evals/ (Kinniment et al., 2023).
Sociotechnical Systems.
The concept of sociotechnical originates from labor studies in coal mining (Trist & Bamforth, 1951; Ropohl, 1999), where the technical components refer to optimizing components of machines themselves, such as data in AI systems, sampling methods, training algorithms, etc. (Weidinger et al., 2023b), while the societal components pertain to the complementary interaction between humans and machines as a social whole. Its core ideas are encapsulated in two principles  [^4] https://open.ncl.ac.uk/theories/9/socio-technical-theory/: Firstly, there is an inherent interaction and mutual influence between social and technical components in defining the success of a system. Secondly, there needs to be a certain degree of fit between social and technical aspects to achieve their joint optimization. Fundamentally, this implies equal consideration of technological and human factors throughout the entire process of societal technological design or redesign (Emery, 1980; Trist, 1981; Emery, 1993).
Lazar & Nelson (2023); Shelby et al. (2023) have revealed that considering AI systems solely from a technical component perspective is insufficient to ensure their safety. It is also necessary to consider the context of human society to ensure genuine safety in the practical deployment of AI systems. Existing research on alignment in sociotechnical systems primarily focuses on the evaluation of societal components. Selbst et al. (2019) identifies five challenges for AI technologies in sociotechnical systems, proposing a shift in design focus towards processes and considering social participants. Weidinger et al. (2023b) advocates for a three-tier evaluation framework to assess AI risks, emphasizing the need for both humans and machines in technology. They also propose methods to bridge safety assessment gaps. Dean et al. (2021) highlights the importance of considering dimensions such as optimization, consensus, value, and failure in AI research, exemplified by drones.
@@ 4 Background and Overview
In this section, we formulate the interactive dynamics among multiple AI and human populations coexisting in a societal context as a multi-player Markov game, referred to as Sociotechnical Game (SG). The game is assumed to take place in an infinite-horizon discounted setting. The complexity of this game can be extended by increasing the number of AI and humans, the frequency of interactions, and introducing elements such as incomplete information and imperfect information. Specific details will be adjusted based on the real-world scenarios and issues discussed later in the text. Here, we present a general formulation.
Formally, the SG is defined as a tuple $\mathcal{G}=(\mathcal{N}^{p},\mathcal{S}^{p},\mathcal{A}^{p},R^{p},\rho^{p},\gamma^{p},\mathbb{P})$ with the set of players $\mathcal{N}^{p}$, the state space $\mathcal{S}^{p}$, the joint action space $\mathcal{A}^{p}$, the reward function $R^{p}$, the initial state distribution $\rho^{p}$, the discount factor $\gamma^{p}$ and the transition probability function $\mathbb{P}$. We use $\mathcal{H}$ and $\mathcal{I}$ to represent human and AI, respectively. $p$ denotes a player. Given that the scenarios considered in this paper involve both human and AI, it is established that $p\in\{\mathcal{H},\mathcal{I}\}$. We use $n$ to denote the total number of players, where $n:=\left|\mathcal{H}^{p}\cup\mathcal{I}^{p}\right|$, In the more detailed definitions provided below, to ensure precision in the formulation, we will distinguish basic symbols in various contexts by incorporating superscripts or subscripts.
4.1 Basic Concepts in Human-AI Interaction
Policies. policy $\pi^{p}_{k}$ for the $k$-th player is a function mapping a given state to a distribution over available actions:
$\pi^{p}_{k}:\mathcal{S}^{p}\ni s\mapsto\pi^{p}_{k}(\cdot\mid s)\in\Delta\left(\mathcal{A}^{p}_{k}\right),$ (1)
where $k\in\mathbb{N}$ and $0<k\leq n$. Policies delineate how AI or human make decisions based on their current state, real-time observations, and the prevailing societal context. We let $\Pi^{p}:\mathcal{S}^{p}\rightarrow\Delta(\mathcal{A}^{p})$ denote the policy space for each player.
Value function. We use $V_{s}(\pi^{p}_{k},\pi^{p}_{-k})$ to denote the value function:
$V_{s}(\pi^{p}_{k},\pi^{p}_{-k}):\mathcal{S}^{p}\ni s,\Pi^{p}\ni{\pi}^{p}_{k}\mapsto\mathbb{R},$ (2)
where $-k$ denotes the set of player indices excluding the index corresponding to the $k$-th player. The value function is defined as the expected cumulative discounted reward at state $s\in\mathcal{S}^{p}$ under the joint policy $({\pi}^{p}_{k},{\pi}^{p}_{-k})$:
$V_{s}\left(\pi^{p}_{k},\pi^{p}_{-k}\right):=\mathbb{E}_{\left(\pi^{p}_{k},\pi^{p}_{-k}\right)}\left[\sum_{t=0}^{\infty}\gamma^{t}R^{p}\left(s_{t},{a}^{k}_{t},a^{-k}_{t}\right)\mid s_{0}=s\right].\footnotesize$ (3)
The reward function $R^{p}$ encapsulates the short-term value of a particular strategy at a given moment. It provides a quantification of the immediate benefits or drawbacks associated with choosing a specific course of action. The initial state is sampled from the initial state distribution $\rho^{p}$. Therefore, the value function is denoted as:
$V_{\rho^{p}}\left(\pi^{p}_{k},\pi^{p}_{-k}\right):=\mathbb{E}_{s\sim\rho^{p}}\left[V_{s}\left(\pi^{p}_{k},\pi^{p}_{-k}\right)\right].$ (4)
The value function captures the enduring values of AI or humans, guiding policy selection in the current state. Estimating and controlling it is crucial for aligning AI with human values, enabling systematic evaluation and management of the long-term implications of current choices. This alignment fosters synergy between AI and human values.
4.2 Overview
Figure 2: The figure illustrates how IC tackles ICSAP based on three classic game-theoretic problems. Figure (a) illustrates the case of IC through mechanism design. The left side of the figure demonstrates a sociotechnical gap between agents considering only technical components and the values of real humans. On the right side, it shows that by designing corresponding mechanisms according to different needs, we can adjust the values of agents, aligning their utility with human requirements under IC conditions, thus achieving alignment in sociotechnical systems. For the left half of (b) and (c), both depict a sociotechnical gap between humans and a single agent. The right half of (b) demonstrates humans designing contracts that satisfy IC conditions based on specific needs, thereby adjusting the values of the agent through the contract. The right half of (c) illustrates a scenario of Bayesian persuasion where humans design information satisfying IC conditions according to their own needs, allowing agents to choose actions maximizing human demands without compromising their own values, thus solving ICSAP.
We consider that AI system agents, which take into account both human factors and only technical components during the development process, collectively form a sociotechnical scenario. We assume that these AI agents exhibit self-interested characteristics and will seek to maximize the value or objectives generated during their training process. In this scenario, we aim to demonstrate that IC can simultaneously address the technical components and societal components of AI systems, enabling AI to achieve consensus with human society in different contexts by adjusting different incentive conditions.
@@ 4.2 Overview
In the following sections, we will demonstrate three classic game problems as applying media of IC properties: Mechanism Design in section 5, Contract Theory in section 6, and Bayesian Persuasion in section 7. An overview of these approaches is depicted in Figure 2.
5 Mechanism Design
In this section, we will discuss how to implement the mechanism design method to apply IC for solving ICSAP.
5.1 Background
@@
Mechanism Design theory deals with private information games where individual types and values are unknown to the designer (Nisan & Ronen, 1999). It typically promotes heterogeneous value agents to reveal their private information and reach equilibrium at desired outcomes by constructing an efficient social structure for incentives (Dafoe et al., 2020; Ji et al., 2023).
@@
In mechanism design, IC is a fundamental constraint, alongside individual rationality, that restricts the possible mechanisms and social functions. However, the revelation principle (Dasgupta et al., 1979) shows that IC doesn’t limit our ability but simplifies strategic behaviors in rule design. It states that every Bayesian-Nash implementable social choice function can be achieved with incentive compatibility, treating IC as a “free lunch” scenario and allowing focus within this context.
@@
Due to the generalized definition and objectives of mechanism design, it finds numerous applications in social choice theory (Gibbard, 1973; Satterthwaite, 1975), voting theory (Dasgupta & Maskin, 2020), stable matching (Gale & Shapley, 1962), and auction theory (Myerson, 1981; Clarke, 1971). For example, Heidari et al. (2018); Huang et al. (2019); Weidinger et al. (2023a) studied and analyzed the impact of the Veil of Ignorance mechanism (Rawls, 1971) on social fairness and found that it promotes societal governance. Sinha & Anastasopoulos (2015); Zheng et al. (2020, 2022) ensure the maximization of social welfare and fairness through algorithmic learning of tax mechanisms.
@@
Among them, the mechanism design has been most widely applied in the auction field. For example, the second-price auction (Vickrey, 1961) is one of the simplest IC mechanisms. In a single-item environment, under the rule where the highest bidder pays the second-highest price, the weakly dominant strategy for bidders is to honestly reveal their valuation. In multi-item scenarios, achieving IC and maximizing social welfare generally rely on the Vickrey-Clarke-Groves (VCG) mechanism (Clarke, 1971). This mechanism aligns bidder utility maximization with social welfare maximization by initially paying each bidder the sum of the others’ valuations, and then using a payment (utility) function based solely on the other bids to ensure IC. By setting the payment function to collect payments equal to the maximum social welfare when the bidder is absent, the designer ensures no net payment is needed, thus accounting for the externalities generated by the bidders.
5.2 Prospects
The general objective of mechanism design is the social choice function, which is defined as the following formula: $f:\bigtimes_{i=1}^{n}\Theta^{p_{i}}\rightarrow\bigtimes_{i=1}^{n}\mathcal{S}^{p_{i}}$, where $\Theta^{p}$ denotes the type of the agent. In the auction setting, type refers to the value of the good to the bidders. For AI agent, it can be viewed as the parameters of a neural network or an indicator of its maximization goal, such as utility or welfare. Social value, within this context, can be described in the form of preference over a set of choice functions $\mathcal{F}_{f}$. Let $F(\cdot):\mathcal{F}_{f}\rightarrow\mathbb{R}$ denote this preference. As an example, we can define the value of fairness for the cake-cutting problem in the following statement:
$\forall f_{1},f_{2},\;\;\max_{i,j}|f_{1,i}-f_{1,j}|$ $\leq\max_{i,j}|f_{2,i}-f_{2,j}|$ (5)
$\Rightarrow F(f_{1})$ $\leq F(f_{2}),$
where $f_{1,i}$ represent the share of cake i-th agent receive and $F$ represent the fairness. All the agents have the same type of greed for cake share. In this framework, the alignment of heterogeneous agents with certain values can be expressed as finding the appropriate mechanism to implement a social choice function that maximizes a targeted social value.
In the auction setting, mechanism design can be separated into two parts, which is the same in our problem setting:
$\text{Allocation rules}\;g:\bigtimes_{i=1}^{n}\mathcal{S}^{p_{i}}\times\bigtimes_{i=1}^{n}\Pi^{p_{i}}\rightarrow\bigtimes_{i=1}^{n}\mathcal{S}^{p_{i}},\vspace{-10pt}$ (6)
$\text{Payment rules}\;c:\bigtimes_{i=1}^{n}\Pi^{p_{i}}\rightarrow\mathbb{R}^{n}.\vspace{-5pt}$ (7)
Allocation rules can be treated as a function that implicitly changes the transition probability function. Payment rules, on the other hand, directly modify the agent’s ultimate utility and are additional degrees of freedom in design to ensure IC property. The main difference between the auction and our problem setting is that the allocation rules are state-dependent, to be used for multiple periods. After placing agents under these rules, their utilities are:
$u_{g,c,k}^{p}(s_{0},\pi^{p}):=\mathbb{E}_{g(\pi_{k}^{p},\pi_{-k}^{p})}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{k}^{p}|s_{0}\right]+c(\pi_{k}^{p},\pi_{-k}^{p}).$ (8)
By assuming the type of the agent is its reward function and an indicator for utility maximization, we have the agent’s optimization problem:
$\max_{\pi_{k}^{p}}\mathbb{E}_{s\sim\rho^{p}}\left[u_{g,c,k}^{p}(s,\pi^{p})\right]\!=\!\max_{\pi_{k}^{p}}U_{g,c,k}^{p}(\pi^{p}),\;\;\forall k\!\in\![n].$ (9)
Using the revelation principle, we can further incorporate the calculation of optimal policy into a part of the mechanism. Denoting the optimal policy as a function of type $\pi_{k}^{p*}(R_{k}^{p})$, the IC condition refers to:
$U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p}),\pi_{-k}^{p*})\geq U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p^{\prime}}),\pi_{-k}^{p*}),\;\;\forall R_{k}^{p^{\prime}}.$ (10)
In order to generalize mechanism design to our ICSAP problems, we require a comprehensive framework for conducting design. One potential option is automated mechanism design (Conitzer & Sandholm, 2002). This theory suggests generating optimal rules and mechanisms for specific problems through a set of procedures. By integrating deep learning (Shen et al., 2018), automated mechanism design has been employed to resolve optimal auction (Dütting et al., 2017). This approach incorporates the IC condition into the optimization objective by introducing regret, thereby achieving an $\epsilon$-IC property.
Therefore, we combine the social value and the penalty on the deviation of IC to generate the loss function for automated mechanism design:
$loss=-F(g,c)+\alpha\sum_{k=1}^{n}rgt_{k}(R_{k}^{p},R_{-k}^{p}),\vspace{-5pt}$ (11)
where $\alpha$ is a scaling parameter and $F$ is social value that takes $(g,c)$ as a representation of social choice function. $rgt_{k}$ function is defined as the difference between the optimal value and the ultimate utility under IC strategy:
$rgt_{k}(R_{k}^{p},R_{-k}^{p})=\max_{R_{k}^{p}}U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p}))-U_{g,c,k}^{p}(\pi_{k}^{p*}(R_{k}^{p})).$ (12)
A certain procedure will be taken to solve this loss minimization problem and come out with the optimal rules.
6 Contract Theory
In this section, we will discuss how to implement the contract theory method to apply IC for solving ICSAP.
6.1 Background

### v3

# 2State Key Laboratory of General Artificial Intelligence, BIGAI

@@ Abstract
Keywords:
incentive compatibility collaborative intelligence AI alignment sociotechnical systems.
@@ 1 Introduction
The rapid development of artificial intelligence (AI) has had a significant impact on human society [48, 58, 70, 67, 80], from robots entering human production and living environments [50, 25] to large language models (LLMs) capable of complex natural language interactions [81, 7] and reasoning ability [73, 72]. The problem will be much more significant for Artificial General Intelligence (AGI). Therefore, an increasing number of people believe that as AI capabilities improve, AI systems will become integrated into human society in the future and be deployed in increasingly complex scenarios [29, 24]. Conversely, the powerful capabilities of AI systems have raised concerns about their safety [10, 58], especially considering their behavioral motivations  [^1] https://www.scai.gov.sg/scai-question-6/, alignment science  [^2] https://www.anthropic.com/news/core-views-on-ai-safety and how they align with human values and intentions [40]. This is recognized as the “AI Alignment” problem.
@@
Substantial progress has been made in addressing AI alignment issues, especially in the forward alignment process [40], which enables AI systems to have alignment capabilities 2. The methods for this process can mainly be divided into two categories. The first category involves learning from feedback [13, 4, 56], and there have been some significant research topics, including preference modeling [76], policy learning [37], and scalable oversight [12, 38, 4, 8]. The second category focuses on resolving distributional shift [21, 52] in learning, with notable subproblems including algorithmic interventions [68, 44, 47], adversarial training [30, 59], and cooperative training [16, 15, 43, 22, 60].
@@
However, these methods only consider the given alignment objectives, focusing solely on technical components such as dataset, architecture, and training algorithms, etc. [75], overlooking the fact that AI systems are sociotechnical systems [64]. Some studies have indicated that relying solely on technical means will result in a sociotechnical gap between the model’s development context and its actual deployment context [64, 45, 79], which is also detrimental to further social governance. Such examples are not uncommon in daily life. ChatGPT, trained on internet data and fine-tuned through RLHF [56], requires prompt engineering for adaptation to individual needs, highlighting unaddressed challenges in existing alignment techniques. Additionally, for sociotechnical systems, existing research is more concerned with only societal components like governance and evaluation methods [20, 75]. Thus, currently, there is a lack of means to simultaneously consider both technical and societal components, enabling AI systems themselves to maintain consensus with human society.
@@
Incentive Compatibility (IC) [36], derived from game theory, suggests that participants only need to pursue their true interests to reach optimal outcomes [62]. This concept leverages self-interested behavior, aligning actions with the game designer’s goals [31]. With IC, each agent can maintain private goal information acquired during pretraining. Only by reconstructing different environments and rules,agents can optimize their own objectives to achieve outcomes that meet the needs of human society in different contexts. Therefore, we believe that exploring the IC property for AI alignment problems in sociotechnical systems is a highly worthwhile research endeavor.
@@
In this paper, we separate a new subproblem from AI alignment problems in sociotechnical systems, called Incentive Compatibility Sociotechnical Alignment Problem (ICSAP), and based on this, we propose our main position: {mdframed}[backgroundcolor=gray!10, roundcorner=10pt] Achieving incentive compatibility can simultaneously consider both technical and societal components in the forward alignment phase, enabling AI systems to keep consensus with human societies in different contexts.
@@ 2 Motivation and Opportunity: A Brief Example
In this section, we will use a very simple example to demonstrate how IC works in addressing AI alignment issues in sociotechnical systems through mechanism design, which will be illustrated specifically in Section 3.1.
@@
Consider a classic divide and choose problem: the two-player cake cutting [66]. In this example (see Figure 1), two self-interested agents aim to maximize their cake share, while the human seeks an equal division. If either agent cuts the cake, they’ll take the whole. To align individual interests with the human’s goal, a simple mechanism is proposed: the cutter chooses second. This constraint ensures the agent’s pursuit of self-interest coincides with the host’s objective, achieving Alignment. The mechanism’s IC conditions facilitate consensus on an equal distribution and maximization of cake consumption.
If we consider the agent as an AI system and its desire to eat the most cake as the objective imparted by the technical component of training, we only need to use automated methods to search for corresponding mechanisms with IC properties as rules based on different real-world requirements to effectively solve ICSAP. Of course, hosts can also have different contextual needs, and they may not necessarily be self-interested. Here, we just provide a possible scenario to illustrate our point.
@@ 3 Background and Overview
In the following sections, we will demonstrate three classic game problems by applying media of IC properties: Mechanism Design in Section 3.1, Contract Theory in Section 3.2, and Bayesian Persuasion in Section 3.3. An overview of these approaches is depicted in Figure 2.
3.1 Mechanism Design
Figure 2: The figure illustrates how IC tackles ICSAP based on three classic game-theoretic problems.
@@ 3.1 Mechanism Design
Mechanism Design theory deals with private information games where individual types and values are unknown to the designer [54]. It typically promotes heterogeneous value agents to reveal their private information and reach equilibrium at desired outcomes by constructing an efficient social structure for incentives [16, 40].
@@
In mechanism design, IC is a fundamental constraint, alongside individual rationality, that restricts the possible mechanisms and social functions. However, the revelation principle [17] shows that IC doesn’t limit our ability but simplifies strategic behaviors in rule design. It states that every Bayesian-Nash implementable social choice function can be achieved with incentive compatibility, treating IC as a “free lunch” scenario and allowing focus within this context.
@@
Due to the generalized definition and objectives of mechanism design, it finds numerous applications in social choice theory [28, 63], voting theory [18], stable matching [26], and auction theory [51, 14, 71]. For example, [33, 35, 74] studied and analyzed the impact of the Veil of Ignorance mechanism [61] on social fairness and found that it promotes societal governance. [65, 82, 83] ensure the maximization of social welfare and fairness through algorithmic learning of tax mechanisms.
@@
Among them, the mechanism design has been most widely applied in the auction field. For example, the second-price auction [69] is one of the simplest IC mechanisms. In a single-item environment, under the rule where the highest bidder pays the second-highest price, the weakly dominant strategy for bidders is to honestly reveal their valuation. In multi-item scenarios, achieving IC and maximizing social welfare generally rely on the Vickrey-Clarke-Groves (VCG) mechanism [14]. This mechanism aligns bidder utility maximization with social welfare maximization by initially paying each bidder the sum of the others’ valuations, and then using a payment (utility) function based solely on the other bids to ensure IC. By setting the payment function to collect payments equal to the maximum social welfare when the bidder is absent, the designer ensures no net payment is needed, thus accounting for the externalities generated by the bidders.
Recent work [49, 55] has constructed environments that encourage people to compete or cooperate through mechanism design. We can similarly apply this approach to AI governance in order to bridge the sociotechnical gap between humans and AI.
The subfigure (a) in Figure 2 illustrates the case of IC through mechanism design. The left side of the figure demonstrates a sociotechnical gap between agents considering only technical components and the values of real humans. On the right side, it shows that by designing corresponding mechanisms according to different needs, we can adjust the values of agents, aligning their utility with human requirements under IC conditions, thus achieving alignment in sociotechnical systems.
3.2 Contract Theory